OpenStudy (anonymous):

a. Find the area of a square garden that is 400/ B2 ft on a side. b. Find the perimeter. HELLLLLLLLP!

OpenStudy (anonymous):

400 over b2

OpenStudy (agent0smith):

Area \[A = (\frac{ 400 }{ B2 })^{2} = \frac{ 400^{2} }{ 4 B ^{2} } = \frac{ 40000 }{ B ^{2} }ft ^{2}\]

OpenStudy (agent0smith):

where?

OpenStudy (agent0smith):

Unless she meant b^2?

OpenStudy (agent0smith):

If it's b^2 and not b*2, \[A = (\frac{ 400 }{ B^2 })^{2} = \frac{ 160000 }{ B^4 } ft ^{2}\]

OpenStudy (agent0smith):

Perimeter would be all four sides added together: \[P = 4 \times \frac{ 400 }{ B ^{2} } =\frac{ 1600 }{ B ^{2} } ft\]